2. A company purchased machinery that cost 510,000. It is estimated that the machine will be operated for 100,000 hours over its useful life and have a residual value of 10,000.






Required:






A) what is the rate of depreciation per hour?






B) journalize the entry for annual depreciation if the machine had been operated for 22,000 hours?

1
Expert's answer
2022-02-15T13:00:22-0500

a)"\frac{510000-10000}{100000}=\frac{500 000}{100 000}=5"


b)"22000\times5=110 000"

Debit main production credit depreciation - 110 000